OpenStudy (anonymous):

please give me a 8 radical equation

OpenStudy (anonymous):

with solution

OpenStudy (anonymous):

\[\sqrt{x}+\sqrt{x}+\sqrt{x}+\sqrt{x}+\sqrt{x}+\sqrt{x}+\sqrt{x}=7\sqrt{x}\]

OpenStudy (anonymous):

or did you want \[\sqrt[8]{n}=2\]or something? This solution is n = 256. The other one is solved by any x>=0.

jhonyy9 (jhonyy9):

radical equation ? like ?

OpenStudy (anonymous):

like this . . . . . . . . . .

OpenStudy (anonymous):

\[72= 8\]

jhonyy9 (jhonyy9):

sqrt(64) =8

jhonyy9 (jhonyy9):

\[\sqrt{64}=8\]

jhonyy9 (jhonyy9):

ok ?